OpenStudy (anonymous):

if f(x) is a cubic polynomial and given the value of f(0) = 10, f(1) = 20, and f(2) = 30. Determine the value of f(2) + f(-2) ?

OpenStudy (anonymous):

f(x) = ax^3 + bx^2 + cx + d

OpenStudy (anonymous):

according information above, i get equations : d = 10 ... (1) a + b + c = 10 ... (2) 6a + 4b + 2c = 20 ... (3) i stuck how to get the value of a, b , and c

OpenStudy (anonymous):

* 8a + 4b + 2c = 20

OpenStudy (anonymous):

f(2) = 8a + 4b + 2c + 10 f(-2) = -8a + 4b - 2c + 10 ------------------------ (+) f(2) + f(-2) = 8b + 20 hmm... still we have b unknown

OpenStudy (anonymous):

yeah, i agree that

OpenStudy (anonymous):

oh, maybe we can take respectively for x = 2, and x = -2 into the 2 equations

OpenStudy (trojanpoem):

Do you mean pluging 2 , -2 in ?

OpenStudy (trojanpoem):

d = 10 a + b +c = 10 4a + 2b + c = 10 f(x) = ax^3 + bx^2 + cx + d f(-x) = -ax^3 +bx^2 -cx + d f(x) + f(-x) = 2bx^2 + 2d f(x) + f(-x) = 2bx^2 + 20

OpenStudy (anonymous):

yeah, but looks doesnt work :) must be solve for b, first

OpenStudy (trojanpoem):

Yeah, we can notice that from the equation f(x) + f(-x) as if we set x = 2 we will get f(2) + f(-2) = 8b + 20

OpenStudy (loser66):

@tanjung can you take a snapshot for the original problem?

OpenStudy (anonymous):

the problem is in indonesian language... i just trying translate it in english and all information is like aobve

OpenStudy (trojanpoem):

So you forgot the most important tip while translating...@tanjung Send it, Loser66 is Indonesian guru. @Loser66

OpenStudy (loser66):

hahaha.... it is funniest thing I 've ever seen. But, just post it, we may use google translation to figure out.

OpenStudy (anonymous):

really ? :)

OpenStudy (loser66):

Cross our fingers. :)

OpenStudy (anonymous):

problem 10, @Loser66

OpenStudy (trojanpoem):

@Loser66, Damn dude, he wasn't joking it's real indonesian.

OpenStudy (anonymous):

yeah, im indonesian @TrojanPoem :)

OpenStudy (loser66):

Yes, it is. :) https://www.google.com/?gws_rd=ssl#q=google+translate

OpenStudy (trojanpoem):

To me, it looks like he translated it properly.

Parth (parthkohli):

Let\[p(x) = x(x-1)(x-2) + 10(x+1)\]It satisfies all the above conditions.

OpenStudy (anonymous):

waw, how can you got fast.... is it a counter solution ?

Parth (parthkohli):

Also it seems that the answer to your question is not unique.\[p(x) = 2x(x-1)(x-2) + 10(x+1)\]\[p(x) = 3x(x-1)(x-2) + 10(x+1)\]all satisfy your condition but return different results.

OpenStudy (anonymous):

so, the answer is many possible right ?

Parth (parthkohli):

Yes, of course.

OpenStudy (amitrashok):

hey can u guys help me with a question sorry to intrude but this is urgent

OpenStudy (amitrashok):

is that ok

Parth (parthkohli):

If you're given any more information, then we can work on the problem.

OpenStudy (anonymous):

oh, ok i got it... thank you very much @ParthKohli

OpenStudy (anonymous):

but what should i do while the question is instruct to me write the answer with the number 0 - 999 ? :)

OpenStudy (loser66):

To me, I have it in the form ax^3 +bx^2 + cx +d with b <0 Since, from above, we have a = -b /3 If b >0, then a <0 but the end of the cubic is positive infinitive, hence a must be >0, that gives us b <0, if b =0, a =0, we don't have cubic. Hence for b <0, let test b =-1, we have \(\dfrac{1}{3}x^3 -x^2 +\dfrac{32}{3}x +10\) satisfies the condition. if b =-2, then \(\dfrac{2}{3}x^3 -2x^2+\dfrac{34}{3}x +10 \) satisfies the condition also. Hence we have infinite many solution for it. To your grade, I think you just pick one of them, like the first one. and calculate f(-2) = -8, then f(2) + f(-2) = 30 -8 =22

OpenStudy (anonymous):

k